OpenStudy (diyadiya):

Prove that the function f given by f(x)=|x-1|, x ∈ R is not differentiable at x = 1.

OpenStudy (anonymous):

You plot the function and show ...

OpenStudy (diyadiya):

Plot?

OpenStudy (anonymous):

Graph

OpenStudy (diyadiya):

i dont think there's a graph hm and idk to plot it ;/

OpenStudy (diyadiya):

ok wait

OpenStudy (turingtest):

the formula\[|x|=\sqrt{x^2}\] should do the trick, right?

OpenStudy (anonymous):

f'(1) should be undefined.

OpenStudy (turingtest):

right, if we use the formula I posted it the derivative will be undefined at 1

OpenStudy (anonymous):

To the right of x=1 line the graph of f(x) has slope +1. To the left of the x=1 it has slope −1.

OpenStudy (anonymous):

However, this description of the derivative leaves out the value of f'(0).

OpenStudy (anonymous):

Alternative analytical way is to show that right hand derivative is not equal to left hand and hence the proof.

OpenStudy (anonymous):

Turing, we should not need any formula to show this.

OpenStudy (diyadiya):

Alternative is better

OpenStudy (mr.math):

Redefine the function as: \[\left| x-1 \right|= x-1 \text{ for } x\ge 1, \] \[\text{ } =1-x \text{ for } x<1\] \[f'(1^+)=1 \ne f'(1^-)=-1.\]

OpenStudy (turingtest):

Hey, I'm letting you run the show here...

OpenStudy (anonymous):

but I would like you to do it geometrically, since it clears intuition :)

OpenStudy (anonymous):

I am grateful to you, Turing :D

OpenStudy (mr.math):

Since \(f'(1^+)\ne f'(1^-)\), then f(x) is not differentiable at x=1.

OpenStudy (diyadiya):

Ok Thanks MrMath! and Thanks FoolForMath & TuringTest :)

OpenStudy (mr.math):

You're welcome!